# Rise of the forsaken relics: connecting present-day stellar streams and phase-mixed galaxies to the Epoch of Reionization

### v1

@@ 5.1 Reconstructed differential SMFs and UVLFs for each present-day classification
5.2 Reconstructed differential SMFs and UVLFs using the full merger histories of present-day substructures
@@
Figures 5 and 6 (top panels) compare the differential SMFs and UVLFs at $z=7$ for Samples I (red), II (blue), and III (green) to those of the proto-MW/LG (black) and the universal SMFs/UVLFs from FIREboxHR (yellow). As discussed in section 5.1, we show the isolated suite (left) and the paired suite (right) separately. Solid lines show median SMFs/UVLFs from all of the simulations in the respective suite. Shaded regions show the 68% halo-to-halo scatter except for the universal FIREboxHR curve (yellow); here the shading denotes the 68% confidence interval (CI) obtained via bootstrapping (Feldmann et al., 2025). We also compare the ratio of the SMF/UVLF from each sample with respect to the SMF/UVLF of the proto-MW/LG (Figures 5 and 6, bottom panels) to help visualize their shapes as a function of stellar mass/UV magnitude. We compute the ratios for each simulation first and then take the median and 68% CI.
The proto-MW/LG SMFs/UVLFs have shapes similar to the universal SMF/UVLF from FIREboxHR, lending confidence to the assumption that the proto-MW/LG provides a representative sample of the global average SMF/UVLF during reionization. However, the amplitude of the proto-MW/LG SMF/UVLF is lower than that of FIREboxHR, particularly in the isolated suite. This could be attributed to the difference in the calculation of the volume that goes into the denominator while computing the SMFs/UVLFs (see section 5). The stellar masses and UV magnitudes in FIREboxHR are also computed differently than in this work (see Feldmann et al. (2025) for details). Additionally, the FIREboxHR box has a higher mean matter density (exactly matching the mean density of the Universe) across the simulation volume than the MW-like zoomed simulations at high redshifts, which could also lead to this difference in amplitude. This discrepancy is smaller in the case of the paired LG-like suite because these simulations have a higher mean matter density than the isolated MW-like simulations, and the galaxies form stars at earlier times in the paired suite, leading to more star-forming galaxies at high redshifts. We plan to explore this difference further in future work.
@@
The overall shapes of the reconstructed SMFs/UVLFs from the three samples are qualitatively similar to that of the proto-MW/LG. Adding the progenitors of the stellar streams recovers the SMF/UVLF of the proto-MW/LG more accurately, particularly in terms of normalization, up to $\sim 50$%. Adding the progenitors of the phase-mixed galaxies improves the reconstruction even more. Including the disrupted systems helps most in recovering intermediate- and high-mass progenitors that are missed when using only intact low-mass galaxies. Furthermore, the halo-to-halo scatter in samples II and III generally appears lower than sample I, with the exception of bottom right panels in Figures 5 and 6. But there the scatter is from just 3 paired halos and hence may not be robust. This is even more apparent when the low-mass/faint-end slope is estimated (Section 6.1). Adding more disrupted galaxies to our present-day samples leads to more robust recovery of our proto-MW/LG SMFs/UVLFs by potentially reducing the halo-to-halo scatter, which is sensitive to the variations in accretion histories. This is particularly prominent in the case of the isolated suite, since it probes a wider range of accretion histories than the paired suite, leading to more scatter in the case of Sample I.
@@
5.3 Reconstructed differential SMFs and UVLFs from the fossil record of present-day substructures
@@
Figure 9: Adding streams and phase-mixed galaxies improves the accuracy and precision of the recovered slope of the differential SMF at low masses. We compare the low-mass slopes of the differential SMFs at $z=7$ for samples I (magenta), II (blue), and III (green) to both the full SMF of the proto-MW/LG (black) and the universal slope estimate from FIREboxHR (yellow). Errorbars indicate the standard deviation on the recovered slope for each individual halo, while the horizontal lines show the median ($\tilde{\mu}_{\alpha}$) of the bootstrapped sample of mean $\alpha$ values across all the halos, computed as described in Section 6.1 and Appendix C. The shaded regions show the median ($\tilde{\sigma}_{\alpha}$) of the bootstrapped sample of standard deviations in $\alpha$, representing the halo-to-halo scatter. We show ($\tilde{\mu}_{\alpha}$) and ($\tilde{\sigma}_{\alpha}$) in the rightmost points for proto-MW and proto-LG by combining them as well as separately. In the all progenitors scenario (top), sample III consistently reproduces the low-mass-end slope of the proto-MW/LG SMF most accurately ($\alpha$ closest to that of the proto-MW/LG for each individual halo) and precisely (smaller errorbars for each individual halo). The median slope ($\tilde{\mu}_{\alpha}$) for sample III is also closest to that of the proto-MW/LG and have the smallest halo-to-halo scatter ($\tilde{\sigma}_{\alpha}$), followed by sample II and I. In the fossil record scenario (bottom), samples II and III still have consistently smaller uncertainties for each individual halo but sample I is more accurate than II and III at recovering the slope in some of the halos. The conclusions remain similar when MW and LG are separately considered.
@@
6 Shapes of the SMFs and UVLFs during the EoR
6.1 Low-mass/faint-end slopes
@@
The figures show slope estimates using progenitor galaxies with $M_{\star}\geq 10^{5}\,M_{\odot}$ or $M_{UV}\leq-10$ (our fiducial choice). We defer a discussion of the effect of varying these cut-offs to Section 6.2. The proto-MW/LG slopes are representative of the universal slopes. There is an offset of $\approx 0.07$ in the median slope values, but they agree within the median halo-to-halo scatter. This could be due to the difference in mean matter densities of the universe and the MW/LG region as discussed in Section 5.2. We will explore this aspect in future work. Since the proto-MW/LG slopes are representative of the universal values within the uncertainties, we mainly assess our results from the three samples against the proto-MW/LG for quantitative comparisons.
@@
We reach similar conclusions when we consider the effect of adding disrupted galaxies separately to a MW-like environment and a LG-like environment. The halo-to-halo scatter in case of a LG-like environment is smaller compared to that of a MW-like environment, but the results are derived from only three simulations and may not be robust. Additionally, the low-mass-end for sample I and III is steeper than proto-LG because of the R&R paired halo (see Figure 9). Furthermore, the slopes are susceptible to how present-day and high-$z$ galaxies are selected, which we discuss in 6.2 and hence the should be treated with caution.
@@
Another important difference can be discerned between the all progenitors scenario and the fossil record scenario: the latter case results in flatter slopes - by $\sim$0.07 on average across the three samples when both the isolated and paired galaxy suites are combined. As discussed in Section 5.3, this effect is due to combining multiple progenitors into a single one in the fossil record scenario and it gets amplified for samples II and III as they have more massive substructures with multiple progenitors. The level of flattening varies depending on how both the present-day and high-$z$ progenitor galaxies are selected.
@@
6.2 Sensitivity to stellar mass selection criteria
In Section 5.3, we noted that the use of the fossil record seemed to increase the sensitivity of our reconstructions to the selection of galaxies used in the near-far technique. Here we explore this further, by testing how the $z=7$ SMF slope estimates change under several variations of our selection criteria.
6.2.1 Lowering the low-mass cutoff at $z=7$
@@
6.2.2 Increasing the present-day stellar mass cutoff for intact galaxies
@@
6.2.3 Removing morphologically complex and extremely massive present-day substructures
@@
6.2.4 Lowering the high-mass cutoff at the present day
@@
6.2.5 Discussion of sensitivity analysis
@@ 6.3 Redshift evolution
In general, sample III consistently reproduces the most accurate and precise proto-MW/LG slope across the redshifts followed by sample II and I, in the all progenitors scenario. The trend is similar in the fossil record scenario, except in a few cases. The reduction in the halo-to-halo scatter with sample II and III is evident across the range of redshifts shown. In particular, at $z=$ 8 and 9, the large scatter in the slope from the progenitors of the intact galaxies alone is fairly well mitigated upon adding the disrupted galaxies, as shown by sample III. We caution, however, that the $z=8$ and $z=9$ results may be subject to the limited statistical samples of progenitor galaxies captured in our simulations at these redshifts. At $z=$ 9, the faint-end slopes from our reconstructed UVLFs are shallower than the universal one. These estimates are probably poor because of the low halo abundance at these redshifts. Overall, for both the SMFs/UVLFs and for most redshifts, sample III reproduces the proto-MW/LG slope more accurately and also reduces the halo-to-halo scatter, followed by sample II and I. This is true for both the all progenitors and fossil record scenarios using our fiducial selection criteria. However, as discussed in section 6.1, the results can vary depending on our sample selection criteria, both at the present-day and at high redshifts.
7 Reconstructed cosmic UV density
Figure 12: Reconstructed cosmic UV density at $z=7$ using the fossil record. Top: contribution of different progenitor galaxies with UV magnitude brighter than $M_{\mathrm{UV}}$ from the fossil records of the different samples to the universal cosmic UV density calculated from FIREboxHR. Solid lines shows the median across all of the halos; shaded regions enclose the 16th to 84th percentile range for the proto-MW/LG (black) and samples I (red), II (blue), and III (green). The orange dotted line shows JWST’s current observational threshold for unlensed galaxies at $z\sim 7$ (Adams et al., 2024). Bottom: relative contribution of different progenitor galaxies to the total reconstructed cosmic UV density within each sample. Including progenitors of disrupted galaxies reduces the uncertainty and predicts greater contributions from brighter galaxies, more representative of the proto-MW/LG. See Section 7 for further discussion.
@@ 7 Reconstructed cosmic UV density
For each of the three samples in each halo, we compute the reconstructed cosmic UV density and explore their contribution to the total cosmic UV density. As shown in Figure 12 (top panel), the proto-MW/LG (which also includes the progenitors of the present-day main host galaxy), can recover between $\sim$18% and $\sim$45% of the total universal cosmic UV density. We cannot fully recover the total universal cosmic UV density because of the difference in normalization between the UVLFs here and that of FIREboxHR, as previously discussed in section 5. Hence, we mainly compare the contribution from the three samples and the proto-MW/LG. Sample III ($\sim$8–18%) is the most complete, followed by sample II ($\sim$4–16%), and sample I ($\sim$1–8%).
@@
8 Detectability of the present-day substructures
@@
To complete our analysis, we construct UVLFs by imposing the detectability constraints obtained from the internal view scenario using LSST. Figure 4—first shown in Section 5.1—gives the differential UVLFs at $z=7$ reconstructed from different classification groups (solid curves) in comparison to reconstructions from their detectable counterparts alone (dashed curves). The reconstructed UVLFs appear promising, even after imposing the detectability constraints. Although the normalization of the satellite and field galaxy progenitor UVLFs is reduced, their shapes are relatively well-preserved. The progenitor UVLFs reconstructed from the streams remain almost unchanged even with just $\sim$50 % detectability, since the detectable streams are mostly of higher mass and each massive stream typically has numerous reionization-era progenitors (see Section 4.2), spanning much of the high-$z$ distribution in progenitor luminosity. We show the results in the all progenitors scenario by utilizing the full merger histories of the present-day substructures, and not in the fossil record scenario. However, the difference between the shape of the UVLFs with and without detectability constraints does not change significantly in the fossil record scenario.
9 Summary
In this work, we test whether including present-day disrupted galaxies like stellar streams and phase-mixed galaxies in the near-far technique improves the inference of the low-mass/faint-end slope of the SMF/UVLF during the EoR. Using the FIRE-2 cosmological zoom-in baryonic simulations of MW-like (isolated) and LG-like (paired) halos (Section 2), we classify present-day substructures into 4 categories: intact satellite galaxies, field galaxies, stellar streams, and phase-mixed galaxies (Section 3). The first two categories constitute our intact galaxies sample (sample I); the intact galaxies and streams comprise our sample II; adding the phase-mixed galaxies makes our sample III. Using particle tracking, we assign each star particle to its progenitor galaxy at high redshifts. We then assess the contribution of different progenitor populations to the total UV luminosity density (Sections 46). Finally, we estimate the proportion of substructures that will be detected by LSST (for the Milky Way) or Roman (for external galaxies) and the effect of these detection limits on the reconstructions (Section 8). Our results are summarized below.

### v2

@@ V.1 Reconstructed differential SMFs and UVLFs for each present-day classification
V.2 Reconstructed differential SMFs and UVLFs using the full merger histories of present-day substructures
@@
Figures 5 and 6 (top panels) compare the differential SMFs and UVLFs at $z=7$ for Samples I (red), II (blue), and III (green) to those of the proto-MW/LG (black) and the universal SMFs/UVLFs from FIREboxHR (yellow). As discussed in section V.1, we show the isolated suite (left) and the paired suite (right) separately. Solid lines show median SMFs/UVLFs from all of the simulations in the respective suite. Shaded regions show the 68% halo-to-halo scatter except for the universal FIREboxHR curve (yellow); here the shading denotes the 68% confidence interval (CI) obtained via bootstrapping (Feldmann et al., 2025). We also compare the ratio of the SMF/UVLF from each sample with respect to the SMF/UVLF of the proto-MW/LG (Figures 5 and 6, bottom panels) to help visualize their shapes as a function of stellar mass/UV magnitude. We compute the ratios for each simulation first and then take the median and 68% CI.
The proto-MW/LG SMFs/UVLFs have shapes similar to the universal SMF/UVLF from FIREboxHR, lending confidence to the assumption that the proto-MW/LG provides a representative sample of the global average SMF/UVLF during reionization. However, the amplitude of the proto-MW/LG SMF/UVLF is lower than that of FIREboxHR, particularly in the isolated suite. This could be attributed to the difference in the calculation of the volume that goes into the denominator while computing the SMFs/UVLFs (see section V). The stellar masses and UV magnitudes in FIREboxHR are also computed differently than in this work (see Feldmann et al. 2025 for details). Additionally, the FIREboxHR box has a higher mean matter density (exactly matching the mean density of the Universe) across the simulation volume than the MW-like zoomed simulations at high redshifts, which could also lead to this difference in amplitude. This discrepancy is smaller in the case of the paired LG-like suite because these simulations have a higher mean matter density than the isolated MW-like simulations, and the galaxies form stars at earlier times in the paired suite, leading to more star-forming galaxies at high redshifts. We plan to explore this difference further in future work.
@@
The overall shapes of the reconstructed SMFs/UVLFs from the three samples are qualitatively similar to that of the proto-MW/LG. Adding the progenitors of the stellar streams recovers the SMF/UVLF of the proto-MW/LG more accurately, particularly in terms of normalization, up to $\sim 50$%. Adding the progenitors of the phase-mixed galaxies improves the reconstruction even more. Including the disrupted systems helps most in recovering intermediate- and high-mass progenitors that are missed when using only intact low-mass galaxies. Furthermore, the halo-to-halo scatter in samples II and III generally appears lower than sample I, with the exception of bottom right panels in Figures 5 and 6. But there the scatter is from just 3 paired halos and hence may not be robust. This is even more apparent when the low-mass/faint-end slope is estimated (Section VI.1). Adding more disrupted galaxies to our present-day samples leads to more robust recovery of our proto-MW/LG SMFs/UVLFs by potentially reducing the halo-to-halo scatter, which is sensitive to the variations in accretion histories. This is particularly prominent in the case of the isolated suite, since it probes a wider range of accretion histories than the paired suite, leading to more scatter in the case of Sample I.
@@
V.3 Reconstructed differential SMFs and UVLFs from the fossil record of present-day substructures
@@
Figure 9: Adding streams and phase-mixed galaxies improves the accuracy and precision of the recovered slope of the differential SMF at low masses. We compare the low-mass slopes of the differential SMFs at $z=7$ for samples I (magenta), II (blue), and III (green) to both the full SMF of the proto-MW/LG (black) and the universal slope estimate from FIREboxHR (yellow). Errorbars indicate the standard deviation on the recovered slope for each individual halo, while the horizontal lines show the median ($\tilde{\mu}_{\alpha}$) of the bootstrapped sample of mean $\alpha$ values across all the halos, computed as described in Section VI.1 and Appendix C. The shaded regions show the median ($\tilde{\sigma}_{\alpha}$) of the bootstrapped sample of standard deviations in $\alpha$, representing the halo-to-halo scatter. We show ($\tilde{\mu}_{\alpha}$) and ($\tilde{\sigma}_{\alpha}$) in the rightmost points for proto-MW and proto-LG by combining them as well as separately. In the all progenitors scenario (top), sample III consistently reproduces the low-mass-end slope of the proto-MW/LG SMF most accurately ($\alpha$ closest to that of the proto-MW/LG for each individual halo) and precisely (smaller errorbars for each individual halo). The median slope ($\tilde{\mu}_{\alpha}$) for sample III is also closest to that of the proto-MW/LG and have the smallest halo-to-halo scatter ($\tilde{\sigma}_{\alpha}$), followed by sample II and I. In the fossil record scenario (bottom), samples II and III still have consistently smaller uncertainties for each individual halo but sample I is more accurate than II and III at recovering the slope in some of the halos. The conclusions remain similar when MW and LG are separately considered.
@@
VI Shapes of the SMFs and UVLFs during the EoR
VI.1 Low-mass/faint-end slopes
@@
The figures show slope estimates using progenitor galaxies with $M_{\star}\geq 10^{5}\,M_{\odot}$ or $M_{UV}\leq-10$ (our fiducial choice). We defer a discussion of the effect of varying these cut-offs to Section VI.2. The proto-MW/LG slopes are representative of the universal slopes. There is an offset of $\approx 0.07$ in the median slope values, but they agree within the median halo-to-halo scatter. This could be due to the difference in mean matter densities of the universe and the MW/LG region as discussed in Section V.2. We will explore this aspect in future work. Since the proto-MW/LG slopes are representative of the universal values within the uncertainties, we mainly assess our results from the three samples against the proto-MW/LG for quantitative comparisons.
@@
We reach similar conclusions when we consider the effect of adding disrupted galaxies separately to a MW-like environment and a LG-like environment. The halo-to-halo scatter in case of a LG-like environment is smaller compared to that of a MW-like environment, but the results are derived from only three simulations and may not be robust. Additionally, the low-mass-end for sample I and III is steeper than proto-LG because of the R&R paired halo (see Figure 9). Furthermore, the slopes are susceptible to how present-day and high-$z$ galaxies are selected, which we discuss in VI.2 and hence the should be treated with caution.
@@
Another important difference can be discerned between the all progenitors scenario and the fossil record scenario: the latter case results in flatter slopes - by $\sim$0.07 on average across the three samples when both the isolated and paired galaxy suites are combined. As discussed in Section V.3, this effect is due to combining multiple progenitors into a single one in the fossil record scenario and it gets amplified for samples II and III as they have more massive substructures with multiple progenitors. The level of flattening varies depending on how both the present-day and high-$z$ progenitor galaxies are selected.
@@
VI.2 Sensitivity to stellar mass selection criteria
In Section V.3, we noted that the use of the fossil record seemed to increase the sensitivity of our reconstructions to the selection of galaxies used in the near-far technique. Here we explore this further, by testing how the $z=7$ SMF slope estimates change under several variations of our selection criteria.
VI.2.1 Lowering the low-mass cutoff at $z=7$
@@
VI.2.2 Increasing the present-day stellar mass cutoff for intact galaxies
@@
VI.2.3 Removing morphologically complex and extremely massive present-day substructures
@@
VI.2.4 Lowering the high-mass cutoff at the present day
@@
VI.2.5 Discussion of sensitivity analysis
@@ VI.3 Redshift evolution
In general, sample III consistently reproduces the most accurate and precise proto-MW/LG slope across the redshifts followed by sample II and I, in the all progenitors scenario. The trend is similar in the fossil record scenario, except in a few cases. The reduction in the halo-to-halo scatter with sample II and III is evident across the range of redshifts shown. In particular, at $z=$ 8 and 9, the large scatter in the slope from the progenitors of the intact galaxies alone is fairly well mitigated upon adding the disrupted galaxies, as shown by sample III. We caution, however, that the $z=8$ and $z=9$ results may be subject to the limited statistical samples of progenitor galaxies captured in our simulations at these redshifts. At $z=$ 9, the faint-end slopes from our reconstructed UVLFs are shallower than the universal one. These estimates are probably poor because of the low halo abundance at these redshifts. Overall, for both the SMFs/UVLFs and for most redshifts, sample III reproduces the proto-MW/LG slope more accurately and also reduces the halo-to-halo scatter, followed by sample II and I. This is true for both the all progenitors and fossil record scenarios using our fiducial selection criteria. However, as discussed in section VI.1, the results can vary depending on our sample selection criteria, both at the present-day and at high redshifts.
VII Reconstructed cosmic UV density
Figure 12: Reconstructed cosmic UV density at $z=7$ using the fossil record. Top: contribution of different progenitor galaxies with UV magnitude brighter than $M_{\mathrm{UV}}$ from the fossil records of the different samples to the universal cosmic UV density calculated from FIREboxHR. Solid lines shows the median across all of the halos; shaded regions enclose the 16th to 84th percentile range for the proto-MW/LG (black) and samples I (red), II (blue), and III (green). The orange dotted line shows JWST’s current observational threshold for unlensed galaxies at $z\sim 7$ (Adams et al., 2024). Bottom: relative contribution of different progenitor galaxies to the total reconstructed cosmic UV density within each sample. Including progenitors of disrupted galaxies reduces the uncertainty and predicts greater contributions from brighter galaxies, more representative of the proto-MW/LG. See Section VII for further discussion.
@@ VII Reconstructed cosmic UV density
For each of the three samples in each halo, we compute the reconstructed cosmic UV density and explore their contribution to the total cosmic UV density. As shown in Figure 12 (top panel), the proto-MW/LG (which also includes the progenitors of the present-day main host galaxy), can recover between $\sim$18% and $\sim$45% of the total universal cosmic UV density. We cannot fully recover the total universal cosmic UV density because of the difference in normalization between the UVLFs here and that of FIREboxHR, as previously discussed in section V. Hence, we mainly compare the contribution from the three samples and the proto-MW/LG. Sample III ($\sim$8–18%) is the most complete, followed by sample II ($\sim$4–16%), and sample I ($\sim$1–8%).
@@
VIII Detectability of the present-day substructures
@@
To complete our analysis, we construct UVLFs by imposing the detectability constraints obtained from the internal view scenario using LSST. Figure 4—first shown in Section V.1—gives the differential UVLFs at $z=7$ reconstructed from different classification groups (solid curves) in comparison to reconstructions from their detectable counterparts alone (dashed curves). The reconstructed UVLFs appear promising, even after imposing the detectability constraints. Although the normalization of the satellite and field galaxy progenitor UVLFs is reduced, their shapes are relatively well-preserved. The progenitor UVLFs reconstructed from the streams remain almost unchanged even with just $\sim$50 % detectability, since the detectable streams are mostly of higher mass and each massive stream typically has numerous reionization-era progenitors (see Section IV.2), spanning much of the high-$z$ distribution in progenitor luminosity. We show the results in the all progenitors scenario by utilizing the full merger histories of the present-day substructures, and not in the fossil record scenario. However, the difference between the shape of the UVLFs with and without detectability constraints does not change significantly in the fossil record scenario.
IX Summary
In this work, we test whether including present-day disrupted galaxies like stellar streams and phase-mixed galaxies in the near-far technique improves the inference of the low-mass/faint-end slope of the SMF/UVLF during the EoR. Using the FIRE-2 cosmological zoom-in baryonic simulations of MW-like (isolated) and LG-like (paired) halos (Section II), we classify present-day substructures into 4 categories: intact satellite galaxies, field galaxies, stellar streams, and phase-mixed galaxies (Section III). The first two categories constitute our intact galaxies sample (sample I); the intact galaxies and streams comprise our sample II; adding the phase-mixed galaxies makes our sample III. Using particle tracking, we assign each star particle to its progenitor galaxy at high redshifts. We then assess the contribution of different progenitor populations to the total UV luminosity density (Sections IVVI). Finally, we estimate the proportion of substructures that will be detected by LSST (for the Milky Way) or Roman (for external galaxies) and the effect of these detection limits on the reconstructions (Section VIII). Our results are summarized below.